hi please help me with this question
A rectangular swimming pool is 8.0 m long. Each cubic metre of water added to the pool increases the depth by 32 mm.
(a) What is the width of the pool?
Round to one decimal place.

m
(b) How much water is needed to fill the pool to a depth of 2.3 m?
Round to the nearest whole.

m3
thanks

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)


A rectangular swimming pool is 8.0 m long. Each cubic metre of water added to the pool increases the depth by 32 mm.
Volume of water 1 cu.m
length of pool = 8.0 m
height per cu.m of water = 0.032 m
Volume = L * W * h
1 = 8* w * 0.032
w= 1/(8*032)
width of the pool = 1.9 m

(b) How much water is needed to fill the pool to a depth of 2.3 m?
Round to the nearest whole.
V = L * W * h
V= 8* 1.9 *2.3
34.96 ~ 35m^3

Answer by ikleyn(53618)

.
A rectangular swimming pool is 8.0 m long. Each cubic metre of water added to the pool increases the depth by 32 mm.
(a) What is the width of the pool?
Round to one decimal place.

(b) How much water is needed to fill the pool to a depth of 2.3 m?
Round to the nearest whole.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        Calculations in the post by @mananth are incorrect.
       I came to bring a correct solution.


(a)   What is the width of the pool?

Volume of water 1 cu.m
length of pool = 8.0 m
height per cu.m of water = 0.032 m
Volume = L * W * h
1 = 8* w * 0.032
w= 1/(8*0.032)
width of the pool = 3.9 m


(b)   How much water is needed to fill the pool to a depth of 2.3 m?
Round to the nearest whole.

V = L * W * h
V = 8 * 3.9 * 2.3
71.76 ~ 72 m^3

Solved correctly.